# SD2: Self-Distilled Sparse Drafters

### v1

@@ Appendix D Detailed SpecBench results
In this section, we report the detailed SpecBench (Xia et al., 2024) results for draft models fine-tuned with SD2 (Table 2), one-shot pruned models (Table 3), SD2 Qwen-2.5 draft models in the UAG setting (Table 4), and compressed & quantized SD2 Llama-3.2 draft models (Table 5). Figure 8 depicts the theoretical improvement factor for Qwen-2.5 draft models, in which SD2 Qwen-2.5 is found to have the highest improvement factor across all candidate draft models.
@@
Table 5: SpecBench results for compressed and quantized SD2 layer-pruned and fine-grained sparse draft models
Llama-1B Dense fp16 $3.91$ $3.98$ $2.96$ $3.41$ $3.59$ $5.62$ $3.65$
SparseGPT fp16 $3.27$ $3.39$ $2.27$ $2.84$ $2.74$ $5.05$ $3.19$
W8A8 $2.76$ $2.84$ $1.89$ $2.39$ $2.29$ $4.60$ $2.70$
Sparse-Marlin $2.70$ $2.86$ $2.04$ $2.08$ $2.36$ $4.70$ $2.47$
Sparse-Marlin fp16 head $3.05$ $3.23$ $2.24$ $2.38$ $2.69$ $4.96$ $2.83$
Llama–3B Dense fp16 $4.47$ $4.50$ $3.58$ $4.14$ $4.27$ $5.90$ $4.28$
SparseGPT fp16 $3.81$ $3.94$ $2.99$ $3.48$ $3.17$ $5.41$ $3.63$
W8A8 $3.43$ $3.55$ $2.54$ $3.12$ $2.81$ $5.08$ $3.31$
Sparse-Marlin $3.31$ $3.48$ $2.69$ $2.77$ $2.83$ $5.05$ $3.03$
Sparse-Marlin fp16 head $3.63$ $3.80$ $2.96$ $3.07$ $3.11$ $5.32$ $3.36$
Appendix E Detailed OpenLLM Leaderboard V1 results
In Tables 6 and 7, we report the benchmark results for our model’s after fine-tuning and after one-shot pruning, respectively. In Table 8, we report the benchmark suite results for the Qwen draft models used in our UAG experiments.

### v2

@@ Appendix F Detailed SpecBench results
In this section, we report the detailed SpecBench (Xia et al., 2024) results for draft models fine-tuned with SD2 (Table 2), one-shot pruned models (Table 3) and SD2 Qwen-2.5 draft models in the UAG setting (Table 4). Figure 9 depicts the theoretical improvement factor for Qwen-2.5 draft models, in which SD2 Qwen-2.5 is found to have the highest improvement factor across all candidate draft models.
@@
Appendix G Detailed OpenLLM Leaderboard V1 results
In Tables 5 and 6, we report the benchmark results for our model’s after fine-tuning and after one-shot pruning, respectively. In Table 7, we report the benchmark suite results for the Qwen draft models used in our UAG experiments.